OpenStudy (anonymous):

Will fan and medal 15% of Mrs. Landry’s students live in Lafayette, Louisiana. Mrs. Landry plans on selecting 15 out of all her students to send postcards to. Find the probability that Mrs. Landry would select exactly 5 students who live in the Lafayette area to send postcards to out of the 15.

OpenStudy (photon336):

Here is an attempted answer. the probability that a student selected out of her class lives in Louisiana is \[0.15\] \[\frac{ number~of~desired~outcomes }{ total~possible~outcomes } = \frac{ 5 }{ 15 }\] i'm not exactly sure but i'm wondering if we can multiply the two events together. \[(\frac{ 5 }{ 15 })(\frac{ 15 }{ 100 }) = (\frac{ 3 }{ 60 })\]

OpenStudy (photon336):

@Kainui

OpenStudy (anonymous):

thank you very much @Photon336 I just could not find the right formula throughout my whole unit. Much appreciated.

OpenStudy (photon336):

Yeah when dealing with any probability problem it's the number of outcomes you're looking for over the total number of outcomes.